OpenStudy (anonymous):

s(x)=x-sinx 0≤x≤4pi a) find the intervals of increase or decrease b) find the local max and min values c) find the intervals of concavity and the inflection points

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

take the derivative, then check to see where it is positive and where it is negative not to hard in this case

OpenStudy (anonymous):

since \(s'(x)=1-\cos(x)\) and since \(|\cos(x)|\leq 1\) this sucker is never negative

OpenStudy (luigi0210):

What the Guru said ^